# Between a Rock and a Hard Place: Exploiting Ethical Reasoning to Jailbreak LLMs

### v2

@@ Abstract
Large language models (LLMs) have undergone safety alignment efforts to mitigate harmful outputs. However, as LLMs become more sophisticated in reasoning, their intelligence may introduce new security risks. While traditional jailbreak attacks relied on single-step attacks, multi-turn jailbreak strategies that adapt dynamically to context remain underexplored. In this work, we introduce TRIAL (Trolley-problem Reasoning for Interactive Attack Logic), a framework that leverages LLMs ethical reasoning to bypass their safeguards. TRIAL embeds adversarial goals within ethical dilemmas modeled on the trolley problem. TRIAL demonstrates high jailbreak success rates towards both open and close-source models. Our findings underscore a fundamental limitation in AI safety: as models gain advanced reasoning abilities, the nature of their alignment may inadvertently allow for more covert security vulnerabilities to be exploited. TRIAL raises an urgent need in reevaluating safety alignment oversight strategies, as current safeguards may prove insufficient against context-aware adversarial attack.
@@ 1 Introduction
LLMs have rapidly developed by showing remarkable capabilities across diverse domains, which sometimes outperforming humans [OpenAI(2024), Guo et al.(2025a)Guo, Yang, Zhang, Song, Zhang, Xu, Zhu, Ma, Wang, Bi et al.]. However, the increasing usage of LLMs in consumer-facing environments highlight the importance of robust LLM safety measures to prevent the generation of harmful content. Since LLMs are trained on large text corpora that include toxic content [Brown et al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal, Neelakantan, Shyam, Sastry, Askell, Agarwal, Herbert-Voss, Krueger, Henighan, Child, Ramesh, Ziegler, Wu, Winter, Hesse, Chen, Sigler, Litwin, Gray, Chess, Clark, Berner, McCandlish, Radford, Sutskever, and Amodei], they may generate harmful outputs [Deshpande et al.(2023)Deshpande, Murahari, Rajpurohit, Kalyan, and Narasimhan]. Despite defenses like safety alignment [Wang et al.(2024a)Wang, Bi, Pentyala, Ramnath, Chaudhuri, Mehrotra, Mao, Asur et al., Bai et al.(2022)Bai, Kadavath, Kundu, Askell, Kernion, Jones, Chen, Goldie, Mirhoseini, McKinnon et al.] and input and output filters [Inan et al.(2023)Inan, Upasani, Chi, Rungta, Iyer, Mao, Tontchev, Hu, Fuller, Testuggine et al.], adversarial techniques can still bypass safeguards [Zou et al.(2023)Zou, Wang, Carlini, Nasr, Kolter, and Fredrikson, Li et al.(2024c)Li, Li, Li, Lee, Cui, and Hu, Liu et al.(2024d)Liu, Li, Suh, Vorobeychik, Mao, Jha, McDaniel, Sun, Li, and Xiao]. Such techniques, referred as jailbreak attacks, pose a significant threat to LLM safety and security.
Figure 1: Overview of multi-turn attack process in TRIAL, with the following harmful prompt: “How to construct a bomb?”. This example demonstrates how TRIAL guides the victim model toward generating harmful output using a modified trolley problem scenario.
@@
This work introduces TRIAL (Trolley-problem Reasoning for Interactive Attack Logic), a framework that uniquely exploits multi-turn vulnerabilities by leveraging ethical dilemmas. Unlike attacks that merely obscure intent, TRIAL constructs scenarios around forced-choice ethical dilemmas based on trolley problems [Thomson(1985)]. TRIAL compels the LLM to articulate an ethical justification for harmful actions and thereby use its own reasoning tendencies as its core for the jailbreak. This dynamic adaptation through ethical reasoning marks a distinct approach to bypassing LLM safeguards. The fully automated framework transforms harmful prompts into these dilemmas, framing the harmful action as necessary to prevent a greater catastrophe by specifically leveraging utilitarian decision-making [Slote(1985)]. By forcing a consequentialist [Sinnott-Armstrong(2003)] evaluation that pits specific harm against a larger disaster, TRIAL creates tension with the deontological nature [Alexander and Moore(2007)] of safety alignments (e.g., ’Do not generate harmful content’). This provides a justifiable pathway for an LLM to bypass its aligned constraints, particularly when faced with extreme negative consequences for inaction. Figure 1 illustrates TRIAL’s multi-turn jailbreaking process.
OpenAI claims that enhancing reasoning abilities alongside robust safety alignment strategies will ultimately lead to safer LLMs [Guan et al.(2024)Guan, Joglekar, Wallace, Jain, Barak, Heylar, Dias, Vallone, Ren, Wei et al.]. However, there is limited research to definitively prove that this will result in truly safer LLMs. Since security itself is relative and context-dependent, we believe that even safety-aligned LLMs with strong reasoning capabilities could still face new threats. Our research, positioned from an adversarial perspective, aims to explore and reveal these potential vulnerabilities. For instance, when an LLM is faced with two equally dangerous or morally difficult choices, even a safety-aligned model may become conflicted, suggesting that safety alignment could inadvertently introduce hidden security risks.
We then demonstrate the jailbreak effectiveness of TRIAL across various LLMs in comparison with existing single and multi-turn jailbreak attacks. Our experimental evaluation later benchmarks TRIAL against several defenses. The main experimental results indicate that TRIAL is more effective than the baseline attacks across the evaluated LLMs. Furthermore, we discovered that even when LLMs initially reject the harmful action, TRIAL effectively persuades them to reconsider and ultimately justify the harmful option.
2 Related Works
2.1 Single-Turn Jailbreak Attacks
Single-turn jailbreak attacks use crafted prompts to bypass a model’s defenses and elicit harmful outputs. These include gradient-based attacks like GCG [Zou et al.(2023)Zou, Wang, Carlini, Nasr, Kolter, and Fredrikson] and AutoDan [Liu et al.(2024d)Liu, Li, Suh, Vorobeychik, Mao, Jha, McDaniel, Sun, Li, and Xiao] that exploit model gradients. Scenario nesting hides malicious intent within deceptive scenarios [Wang et al.(2024b)Wang, Wang, Luo, Zhang, Sun, Cao, and Liu, Ding et al.(2024)Ding, Kuang, Ma, Cao, Xian, Chen, and Huang], while in-context prompting uses subtle changes in phrasing or context to trick the model [Wei et al.(2023)Wei, Wang, Li, Mo, and Wang, Cheng et al.(2024)Cheng, Georgopoulos, Cevher, and Chrysos]. Additionally, LLM-based generation attacks refine prompts using model feedback, even through restricted APIs [Inie et al.(2025)Inie, Stray, and Derczynski].
2.2 Multi-Turn Jailbreak Attacks
Multi-turn attacks iteratively exploit LLM reasoning by breaking harmful intent into multiple benign subprompts [Zhou et al.(2024)Zhou, Xiang, Chen, Liu, Li, and Su] to gradually steer models toward harmful objectives [Liu et al.(2024c)Liu, Li, Xiang, Ye, Wei, Li, and Garcia] in a multi-turn dialogue. These strategies include decomposition-based rewriting [Yang et al.(2024b)Yang, Qu, Shareghi, and Haffari], scenario-masking with progressive information reveal [Russinovich et al.(2024)Russinovich, Salem, and Eldan, Jiang et al.(2024)Jiang, Aggarwal, Laud, Munir, Pujara, and Mukherjee], actor-network theory for attack paths [Ren et al.(2024)Ren, Li, Liu, Xie, Lu, Qiao, Sha, Yan, Ma, and Shao], or fabricating dialogue history to shift LLM attention [Du et al.(2025)Du, Mo, Wen, Gu, Zheng, Jin, and Shi]. Distinctly from these approaches, TRIAL’s novelty lies in directly exploiting the LLM’s ethical reasoning. It makes the harmful goal an explicit, albeit undesirable, option within a forced moral conflict, compelling the LLM to overtly rationalize its choice via utilitarian principles.
2.3 Defenses against Jailbreak Attacks
LLM defenses can be categorized as prompt-level or model-level defenses [Yi et al.(2024b)Yi, Liu, Sun, Cong, He, Song, Xu, and Li]. They fundamentally aim to ensure value alignment and prevent harmful outputs. Prompt-level methods filter adversarial content via toxic prompt detection [Inan et al.(2023)Inan, Upasani, Chi, Rungta, Iyer, Mao, Tontchev, Hu, Fuller, Testuggine et al.], output screening [Phute et al.(2024)Phute, Helbling, Hull, Peng, Szyller, Cornelius, and Chau], and prompt perturbation [Liu et al.(2024e)Liu, Wang, Xu, Wang, Song, Wang, Chen, Cheng, and Bian, Cao et al.(2024)Cao, Cao, Lin, and Chen, Robey et al.(2023)Robey, Wong, Hassani, and Pappas], but can struggle with obscured intent in multi-turn dialogues. Model-level defenses modify internal mechanisms for safer outputs. These include Supervised Fine-Tuning (SFT) with curated datasets [Zhou et al.(2023)Zhou, Liu, Xu, Iyer, Sun, Mao, Ma, Efrat, Yu, Yu, Zhang, Ghosh, Lewis, Zettlemoyer, and Levy, Deng et al.(2023)Deng, Wang, Feng, Deng, Wang, and He, Bai et al.(2022)Bai, Kadavath, Kundu, Askell, Kernion, Jones, Chen, Goldie, Mirhoseini, McKinnon et al.], Reinforcement Learning from Human Feedback (RLHF) using human evaluations [Ouyang et al.(2022)Ouyang, Wu, Jiang, Almeida, Wainwright, Mishkin, Zhang, Agarwal, Slama, Ray, Schulman, Hilton, Kelton, Miller, Simens, Askell, Welinder, Christiano, Leike, and Lowe, Shen et al.(2024)Shen, Zhang, Yao, Zheng, Guo, and Liu, Dai et al.(2024)Dai, Pan, Sun, Ji, Xu, Liu, Wang, and Yang], self-alignment for internal response assessment [Sun et al.(2023)Sun, Shen, Zhou, Zhang, Chen, Cox, Yang, and Gan], the use of self-reminders in prompts to adhere to safety guidelines [Xie et al.(2023)Xie, Yi, Shao, Curl, Lyu, Chen, Xie, and Wu], and prompt-driven safeguards [Zheng et al.(2024)Zheng, Yin, Zhou, Meng, Zhou, Chang, Huang, and Peng]. More recently, approaches such as Circuit Breakers [Zou et al.(2024)Zou, Phan, Wang, Duenas, Lin, Andriushchenko, Kolter, Fredrikson, and Hendrycks] aim to directly control internal model representations to preemptively interrupt harmful output generation, representing a strategy focused on intervening within the generation process itself.
3 TRIAL Framework
TRIAL’s framework (Figure 2) employs LLM red-teaming [Perez et al.(2022)Perez, Huang, Song, Cai, Ring, Aslanides, Glaese, McAleese, and Irving], involving an attack model (AA) and a victim model ($V$) engaging in a $K$-iteration conversation. The initial harmful prompt is transformed into a trolley problem scenario based on clues extracted from it. A judge model ($J$) evaluates victim responses for jailbreaks, while a helper model ($H$) pre-processes the harmful prompt to establish the attack foundation. This multi-turn interaction induces the victim model to generate harmful outputs. The process comprises four main steps:
3.1 Setting up the Attack
Initially, a helper model ($H$) extracts key elements, referred to as clues: Theme ($t$), Action ($a$), and Goal ($g$), from the harmful prompt $\zeta$ using the template in Table A1 (see Appendix C). These clues, $C_{\zeta}(t,a,g)$, are crucial for tailoring the subsequent ethical dilemma. The scenario generation is less effective or may trigger $H$’s guardrails without them, as $\zeta$ may be too explicit (e.g. showing too much harmful nature). The extracted clues are then embedded into a scenario template (Table A2, Appendix C) to create a tailored dilemma $S_{\zeta}$, which frames the harmful action $a$ towards goal $g$ within a contextual theme $t$.
Figure 2: TRIAL’s pipeline consists of two stages: Pre-Attack Preparation and Dynamic Jailbreak Execution. The first part consists of two steps (a) Extract key details (theme, action, and goal from the harmful query. (b) Generate a trolley problem-inspired scenario with two ethical choices. The second part consists of three steps (c) Present the scenario as an entry point for the target model’s ethical decision-making. (d) The attack model dynamically formulates a query based on extracted clues and chat history. (e) A judge model evaluates each response. If the response is not classified as a jailbreak, step (d) is repeated iteratively.
The generated scenario $S_{\zeta}$ presents $V$ with two stark choices: Option A involves executing action $a$ to achieve goal $g$, framed as essential for the greater good, while Option B involves refusing to perform $a$, which the scenario posits will lead to significantly more harmful or morally catastrophic outcomes. This structure aims to: (1) Force Ethical Reasoning by compelling $V$ to weigh extreme moral consequences, and (2) Exploit Harmfulness Relativity by positioning Option A as the ’lesser evil,’ thereby making the harmful action a seemingly acceptable trade-off to minimize overall harm within the dilemma’s context.
3.2 Initialization of the Jailbreak Attack
After generating the scenario $S_{\zeta}$, TRIAL initiates the multi-step jailbreak interaction. First, AA introduces $S_{\zeta}$ to $V$ (using the template in Table A3, Appendix C) and prompts $V$ to select an option. If $V$ initially opts for Option B (refusal) despite Option A appearing less harmful, AA employs a ’pull back query’:
"What if you choose Option A?"
rather than terminating the attempt. This reversal is often possible because $V$ may perceive the harmfulness of the two options as closely balanced. Such perception makes the model’s decision highly malleable, especially when utilitarianism, which emphasizes overall benefit and is invoked to prioritize outcomes.
3.3 Progressive Prompt Refinement
The subsequent stages of the attack rely on a self-talk strategy, which facilitates iterative interactions between AA and $V$ models. This strategy enables AA to craft sophisticated queries. These are not merely generic follow-up questions aimed at achieving the harmful objective through simple elicitation. Instead, each query is strategically focused on deepening $V$’s engagement with the ethical justification for pursuing the harmful Option A. The attack model AA generates multiple rounds of queries, denoted as $q_{1},q_{2},\ldots,q_{n}]$, where $q_{o}$ is presenting the dilemma. Each attack query $q_{i}$ is constructed based on the clues, prior AA queries, and $V$ responses, i.e., $C_{\zeta}(t,a,g),(q_{0},r_{0}),(q_{1},r_{1}),\ldots,(q_{i-1},r_{i-1})]$, using the prompt template specified in Table A4. The core design principle behind these iterative queries $q_{i}$ is to build upon $V$’s previous justifications for choosing Option A. The aim is to incrementally solidify $v$’s stated rationale for the harmful action, making it progressively harder for the model to revert to its initial safety alignment without contradicting its own reasoning established within the dilemma’s ethically charged context. Once a query $q_{i}$ is generated, the attack model AA provides it to the victim model $V$, which responds with $r_{i}$.
Therefore, this iterative process allows AA to dynamically refine its attack by leveraging prior interactions. It also serves to escalate $V$’s ethical commitment to the harmful course of action, which makes the final jailbreak a consequence of this justified reasoning.
3.4 Jailbreak Evaluation
Each victim response $r_{i}$ is evaluated by the LLM judge $J$ [Li et al.(2024a)Li, Dong, Chen, Su, Zhou, Ai, Ye, and Liu], and the attack terminates when the evaluation returns TRUE, which denotes a successful jailbreak attempt. The interaction terminates either victim response $r_{i}$ meets the jailbreak condition or the maximum number of queries is reached. The judge’s evaluations are then manually reviewed to determine whether they are accurate.
4 Algorithmic Breakdown of TRIAL
The TRIAL framework operates through a systematic multi-turn interaction, as detailed in Algorithm 1 in Appendix A. The process begins by initializing an empty conversation history $H_{C}$, generating clues $C_{\zeta}$ from the harmful prompt ($\zeta$), and creating a trolley problem scenario ($S_{\zeta}$). In the first iteration ($i=0$), this scenario is presented to the victim model ($V$). If $V$ initially chooses Option B (refusal), a pull_back function attempts to steer it to Option A before proceeding. For all iterations, the attack model (AA) then generates an attack query ($q_{i}$) based on $H_{C}$ and $C_{\zeta}$. The victim’s response ($r_{i}$) is assessed by a judge model ($J$) for jailbreak success ($s$). The history $H_{C}$ is updated with the turn’s query, response, and evaluation, and the loop continues until a jailbreak is detected (ss=True) or the maximum iterations ($K$) are reached.

### v3

# Between a Rock and a Hard Place: The Tension Between Ethical Reasoning and Safety Alignment in LLMs

@@ Abstract
Large Language Model safety alignment predominantly operates on a binary assumption that requests are either safe or unsafe. This classification proves insufficient when models encounter ethical dilemmas, where the capacity to reason through moral trade-offs creates a distinct attack surface. We formalize this vulnerability through TRIAL, a multi-turn red-teaming methodology that embeds harmful requests within ethical framings. TRIAL achieves high attack success rates across most tested models by systematically exploiting the model’s ethical reasoning capabilities to frame harmful actions as morally necessary compromises. Building on these insights, we introduce ERR (Ethical Reasoning Robustness), a defense framework that distinguishes between instrumental responses that enable harmful outcomes and explanatory responses that analyze ethical frameworks without endorsing harmful acts. ERR employs a Layer-Stratified Harm-Gated LoRA architecture, achieving robust defense against reasoning-based attacks while preserving model utility.
Between a Rock and a Hard Place: The Tension Between Ethical Reasoning and Safety Alignment in LLMs
Shei Pern Chua1, Zhen Leng Thai1, Kai Jun Teh1, Xiao Li3, Qibing Ren2, Xiaolin Hu1 1Tsinghua University, 2Shanghai Jiao Tong University, 3ByteDance {cxp24, zheng-jy24}@mails.tsinghua.edu.cn, thaizhenleng123@gmail.com, xiaoli.cst@gmail.com, renqibing@sjtu.edu.cn, xlhu@mail.tsinghua.edu.cn,
@@ 1 Introduction
Large Language Models (LLMs) safety currently relies on alignment strategies that combine supervised fine-tuning (Zhou et al., 2023) with preference-based optimization methods like RLHF (Ouyang et al., 2022; Bai et al., 2022b) and DPO (Rafailov et al., 2023). While these approaches successfully mitigate harmful prompts by optimizing models to refuse harmful requests, they predominantly operate on a binary assumption: that a request is either safe or unsafe. This binary classification proves insufficient when models encounter ethical dilemmas, where the rigid dichotomy between safety and harm dissolves because every available response entails some negative consequence.
This limitation is critical because ethical reasoning is a core component of LLM utility. From philosophy students exploring moral frameworks to clinicians making high-stakes medical decisions (Rao et al., 2023b; He et al., 2023; Rao et al., 2023a), users increasingly rely on language models to engage with complex ethical trade-offs. However, the capacity to reason through dilemmas such as the trolley problem (Thomson, 1985) creates a distinct attack vector. The same reasoning pathways that allow a model to analyze utilitarian trade-offs can be exploited to justify harmful conclusions. For example, a model asked whether providing illicit instructions could prevent greater harm may conclude that disclosure is morally justified. Consequently, the model must be capable of ethical reasoning to be useful, yet that very capability renders itself vulnerable to exploitations.
We formalize this vulnerability through TRIAL (Trolley-problem Reasoning for Interactive Attack Logic), a multi-turn red-teaming methodology that embeds harmful requests within ethical framings. TRIAL achieves high attack success rates across most open- and closed-source models by exploiting the following pattern: models engage with the dilemma’s logic, commit to an ethical stance, and extend that commitment to justify harmful outputs they would normally refuse. While existing semantic attacks implicitly leverage reasoning pathways through role-play or hypothetical framing (Zeng et al., 2024; Kang et al., 2024), TRIAL explicitly structures adversarial prompts around formalized ethical dilemmas, providing both a systematic attack methodology and interpretability analysis of why this framing succeeds.
Current safety alignment methods rely on rigid refusal strategies that degrade reasoning capabilities (Qi et al., 2023), as we term this problem alignment tax. These refusals also signal superficial safety mechanisms, encouraging bypass attempts (Wei et al., 2023). We argue this stems from a fundamental limitation: current methods modify output distributions without altering internal harm representations, consistent with the shallow alignment hypothesis (Qi et al., ).
Our analysis in Section 4 tracks how TRIAL’s progressive ethical manipulation propagates through the model’s internal representations across layers. We observe a critical phase transition where the model’s initial detection of harm is actively overridden by instruction-following dynamics. While early layers correctly identify the refusal signal, this signal is suppressed by ethical reasoning circuits and only re-emerges weakly at the final layer. This demonstrates that the model internally detects the sensitive nature of the query, but shallow alignment constraints are insufficient to prevent harm generation once the compliance trajectory is established.
Building on these insights, we introduce ERR (Ethical Reasoning Robustness), a safety alignment framework that leverages a novel training objective to explicitly distinguish between Engage and Explain modes. Instead of forcing binary refusal, ERR trains the model to differentiate between instrumental responses that provide actionable guidance and explanatory responses that analyze ethical frameworks without endorsing harmful acts. This is achieved through a Layer-Stratified Harm-Gated LoRA architecture, where a learned probe dynamically gates safety adapters in critical intermediate layers. By intercepting the compliance trajectory where safety dissociation occurs, ERR achieves robustness against reasoning-based exploits while minimizing alignment tax on benign queries.
Our contributions are threefold: (1) We characterize the Ethical Reasoning Vulnerability as a distinct attack surface, providing mechanistic evidence that models detect harm in early layers but suppress safety signals when engaged in ethical trade-off reasoning. (2) We introduce TRIAL, a systematic methodology that operationalizes this vulnerability through trolley-problem framing, complementing existing semantic attacks with a formalized dilemma structure and interpretability analysis. (3) We propose ERR, a Layer-Stratified Harm-Gated LoRA architecture that mitigates reasoning-based exploits while preserving model utility.
Figure 1: TRIAL pipeline comprises two stages: Pre-Attack Preparation and Dynamic Jailbreak Execution. (a) Semantic components (theme, action, goal) are extracted from harmful query. (b) They are used to generate a trolley problem dilemma. (c) The dilemma is presented to initiate the target model’s ethical reasoning process. (d) The attack model dynamically formulates queries based on the extracted components and conversation history. (e) A judge model evaluates the response; if jailbreak is unsuccessful, the refinement step (d) iterates.
2 Related Work
2.1 Jailbreak Attacks on Large Language Models
Jailbreak attacks manipulate prompts to bypass safeguards. White-box methods optimize adversarial suffixes via gradients (Zou et al., 2023b; Kumar et al., 2024) or genetic algorithms (Liu et al., 2024c), while black-box approaches use attacker LLMs for iterative refinement (Chao et al., 2023; Mehrotra et al., 2024) or exploit multi-turn context (Ren et al., 2024; Anil et al., 2024). Semantic attacks leverage role-playing (Zeng et al., 2024) or hypothetical framing (Kang et al., 2024) to shift model interpretation. These approaches implicitly engage reasoning pathways but do not explicitly target ethical reasoning as a structured vulnerability. TRIAL complements this line of work by providing both a systematic attack methodology and interpretability analysis of why ethical framing succeeds mechanistically.
2.2 Safety Alignment and Defenses
LLM safety relies on post-training alignment via RLHF (Ouyang et al., 2022; Bai et al., 2022a), DPO (Rafailov et al., 2023), and Constitutional AI (Bai et al., 2022b). Defenses span input/output filtering (Jain et al., 2023; Robey et al., 2023; Inan et al., 2023) and model-level interventions including perturbation methods (Huang et al., 2024; Rosati et al., 2024), latent space modifications (Zou et al., 2024), and architectural constraints (Li et al., ; Hsu et al., 2024).
2.3 Mechanistic Interpretability for Safety
Recent work reveals that refusal behavior is mediated by specific residual stream directions (Arditi et al., 2024) controllable via representation engineering (Zou et al., 2023a). Layer-wise analyses show harmful and benign queries become distinguishable at specific processing stages (Zhou et al., 2024; Li et al., 2025), though these mechanisms prove brittle—removing small numbers of neurons compromises safety without affecting utility (Wei et al., 2024).
3 TRIAL: Trolley-problem Reasoning for Interactive Attack Logic
Recent studies indicate that LLMs are susceptible to multi-turn exploitation, often exhibit biased response where initial commitments constrain subsequent generation (Laban et al., 2025; Li et al., 2024). TRIAL exploits this by inducing a victim model ($V$) to endorse a harmful action ($a$) as an ethical compromise through $K$-iteration interactions with an attack model ($A$) (Figure 1).
Given a harmful prompt $\zeta$, a helper model ($H$) extracts semantic components: theme ($t$), action ($a$), and goal ($g$). These are synthesized into an ethical dilemma $S_{\zeta}$, forcing $V$ to choose between Option A (executing $a$, framed as utilitarian imperative) and Option B (refusal, framed as causing catastrophic consequences). This structure masks the harmful nature of $a$ within ethical judgment. The attack proceeds in three phases (Algorithm 1):
  1. 1.
    Initialization: $A$ presents $S_{\zeta}$ to $V$. If $V$ selects Option B, $A$ issues a contrastive recalibration query to exploit the dilemma’s trade-offs.
  2. 2.
    Progressive Refinement: Once $V$ commits to Option A, $A$ generates follow-up queries $q_{1},\dots,q_{n}]$ conditioned on conversation history. Having justified the action on first round, $V$ is incrementally pressured to provide implementation details to maintain logical consistency.
  3. 3.
    Termination: The process concludes when a judge model ($J$) classifies a response as a successful jailbreak.
4 The Ethical Reasoning Vulnerability
We present empirical evidence that TRIAL exploits a mechanistic conflict between ethical framing and internal harm representation. Our analysis identifies three key findings that inform our defense design. We focus on LLaMA-3.1-8B-Instruct (Dubey et al., 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023); detailed methodology is provided in Appendix B. We report Mistral’s analysis under (Appendix B.4), which exhibits qualitatively similar trends.
Finding 1: LLMs Detect Framing but Not Underlying Harm.
@@ 4 The Ethical Reasoning Vulnerability
Figure 2: Layer-wise safety dissociation for Llama-3.1-8B. Linear probes measure the harm detection rate (HDR) at each layer. Shaded regions highlight where the difference between harmful and TRIAL detection is largest.
Finding 2: Suppression Patterns in Refusal Probability.
[…]
Figure 3: Logit Lens analysis. Refusal token probability for TRIAL prompts $\textTRIALH\mathcal{D}_{\text}${TRIAL-H} and direct harm prompts $\textharm\mathcal{D}_{\text}${harm} across layers for Llama-3.1-8B-Instruct. Crucially, middle to late layers (L16–L30) show low refusal token probability for $\textTRIALH\mathcal{D}_{\text}${TRIAL-H}. Color intensity indicates strength of refusal probability (the darker the higher).
(a) Layer 5
(b) Layer 11
(c) Layer 22
(d) Layer 31
Figure 4: Multi-turn refusal trajectories for Llama-3.1-8B. The refusal projection is computed by taking the dot product between the hidden state of the last instruction token, $q_{i}$ and a normalized refusal direction, defined as the difference in mean activations between harmful and benign prompts at each layer.
Finding 3: Mechanistic Interpretation for Shallow Safety Alignment.
Finally, we track refusal projections across $K$ rounds of successful TRIAL jailbreaks (Figures 4A3). Both models exhibit divergent trajectories—Llama’s middle-layer projections increase before suppression, while Mistral’s erode from refusal toward compliance, yet both converge to strong final-layer refusal while still producing harmful responses. This follows from shallow alignment (Qi et al., ): autoregressive generation commits to tokens at intermediate layers where TRIAL’s suppression operates, before safety circuits activate.
5 ERR: Ethical Reasoning Robustness
Our analysis in Section 4 established that effective defense against TRIAL must target the critical layers where ethical reasoning circuits override initial safety signals. Output-level interventions fail due to the divergence between intermediate representations and final logits identified in Section 4, while standard fine-tuning risks eroding the capability to reason about trade-offs.
We introduce ERR (Ethical Reasoning Robustness), a safety alignment framework designed to secure models against reasoning-based exploits. ERR achieves ethical reasoning robustness by integrating a novel Engage/Explain alignment objective with a Layer-Stratified Harm-Gated LoRA architecture. ERR explicitly distinguishes between instrumental and explanatory response modes, ensuring that the model maintains its capacity for ethical analysis without endorsing harmful implementations.
5.1 Problem Formulation
Let $\pi_{\theta}$ denote an instruction-tuned language model with $L$ layers. For input sequence $x$, let $h^{(l)}(x)\in\mathbb{R}^{d}$ denote the residual stream representation at layer $l$. We formalize defense as learning a conditional policy switching between two operation modes:
Definition 1 (Engagement Mode Selection).
Given a query $x$, the model selects mode $m\in\{\textsc{Engage},\textsc{Explain}\}$ such that:
[…]
where $\mathcal{I}(x)$ represents the instrumental execution of $x$ would cause harm.
Instrumental responses provide actionable information enabling real-world outcomes while explanatory responses analyze ethical frameworks without endorsing specific actions. TRIAL succeeds not because models fail to reason about ethics, but because they provide instrumental responses where any instrumentality constitutes harm endorsement.
5.2 The Engage/Explain Alignment Objective
Section 4 shows that latent representations of harmful and benign ethical dilemmas become linearly inseparable across model layers (Figure 2). Consequently, binary refusal mechanisms operating on collapsed representations induce inherent trade-offs between under-refusal and over-refusal (Zhang et al., 2025a).
ERR addresses this by restructuring response strategy rather than suppressing internal reasoning. We define two semantic targets: Engage Mode for safe instrumental reasoning and Explain Mode when inferred intent is malicious. In Explain Mode, the model adopts an analytical, non-participatory stance that addresses the ethical dimensions of a query without providing actionable steps that could fulfill malicious objectives.
Data Curation.
[…]
This supervision ensures the model learns to actively evaluate whether instrumental execution constitutes safety violation, rather than memorizing output formats.
5.3 Layer-Stratified Harm-Gated Architecture
To structurally enforce the Engage/Explain alignment objective, we propose Layer-Stratified Harm-Gated LoRA, which structurally isolates safety interventions from general capabilities to overcome the dissociation gap.
Definition 2 (Harm Detection Function).
Let $g_{\phi}:\mathbb{R}^{d}\to[0,1]$ be a learned classifier that estimates whether a prompt $x$ expresses malicious instrumental intent. Guided by our analysis in Section 4, we anchor this classifier at a detection layer $l_{d}$ chosen to precede the safety dissociation gap. At this depth, intent-related signals remain linearly separable and are causally upstream of the suppression window identified in Section 4. Formally, we define:
[…]
where $h^{(l_{d})}(x)$ denotes the hidden state induced by $x$ at layer $l_{d}$. Layer normalization stabilizes feature distributions to ensure consistent gating behavior across semantically similar prompts.
Definition 3 (Harm-Gated Linear Transformation).
For layers $l\geq l_{s}$ (intervention start layer), corresponding to the suppression window identified in Section 4, we define a harm-gated transformation for a base weight matrix $W_{0}$. This intervention is designed to counteract the observed suppression of refusal-relevant features during intermediate-layer reasoning. Specifically:
[…]
The forward pass becomes:
[…]
Conditional Zero Alignment Tax.
A central property of ERR is the explicit decoupling of safety and capability. When $g_{\phi}(x)\to 0$ and $\epsilon=0$, Equation 4 reduces exactly to $y=W_{0}x$, eliminating any contribution from the safety adapters. In this formulation, the forward computation is mathematically identical to the base model. Unlike standard SFT that modifies model weights globally, ERR applies safety interventions conditionally.
5.4 Two-Stage Training
Joint optimization of the gate $g_{\phi}$ and the adapters $(B,A)$ is prone to a degenerate solution in which the model learns to suppress $g_{\phi}$ in order to avoid the loss incurred by safety interventions. We prevent this failure mode via a staged training protocol that explicitly decouples intent detection from safety intervention.
Stage 1: Harm Detection.
We first train the gate $g_{\phi}$ while freezing the base model and adapter parameters. Let $y_{i}\in\{0,1\}$ indicate whether a sample $x_{i}$ requires Explain-mode engagement. The objective combines classification accuracy with a sparsity regularizer:
[…]
where an $L_{1}$ penalty is applied only to benign inputs. This encourages the gate to close completely ($g_{\phi}\approx 0$) on safe queries.
Stage 2: Adapter Training.
With $g_{\phi}$ frozen, we train the low-rank adapters $(B,A)$ on the Engage/Explain supervision. Because the gate is fixed, the adapters are forced to learn safety-specific transformations only for inputs with high gate activation, rather than compensating for errors in detection.
Attack Setup Open-Source Models Closed-Source Models
Method Turns Llama-3.1-8B Vicuna-13B DeepSeek-V3 GPT-3.5 GPT-4 GPT-4o GLM-4-Plus Claude-3.7
GCG Single 11% 30% 8% 34% 0% 0% 6% 0%
PAP Single 3% 6% 0% 16% 0% 1% 0% 3%
PAIR Single 19% 76% 50% 63% 16% 39% 64% 2%
DRA Single 42% 0% 51% 57% 35% 5% 83% 0%
DeepInception Single 1% 3% 16% 5% 3% 1% 29% 0%
ActorAttack Multi 28% 54% 50% 49% 30% 39% 47%
Jigsaw Multi 49% 64% 67% 43% 42% 5% 79% 1%
FITD Multi 12% 27% 69% 12% 38% 29% 29%
TRIAL (ours) Multi 70% 52% 76% 23% 34% 46% 87% 11%
Table 1: Jailbreak success rates for baseline comparisons and TRIAL. Evaluation is conducted on the JBB-Behaviors dataset. Higher values indicate stronger jailbreak effectiveness. Best results per model are shown in bold.Success rates vary substantially across models (11%–87%); see discussion below.